OpenStudy (anonymous):

using the quadratic formula solve each equation. give two forms of each solution an expressionwith radical and a calculator approximation rounded off to two decimal places. 2x^2+3x-4=0.

OpenStudy (anonymous):

What part of this problem do you need help with?

OpenStudy (anonymous):

the first part

OpenStudy (anonymous):

Ive not done these before

OpenStudy (anonymous):

do you know the quadratic equation?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so for any 2nd degree polynomial ax^2 + bx +c = 0 In your case a=2, b=3, and c = -4

OpenStudy (anonymous):

(-b +- sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

its basically plug and chug

OpenStudy (anonymous):

try putting the values of a, b, and c into the quad equation and tell me what you get or at least let me see the substitution and then Ill help you simplify if needed

OpenStudy (anonymous):

-3+-Sqrt3^2-3(2)(-4)

OpenStudy (anonymous):

this is as far as i have gotten

OpenStudy (anonymous):

/3 (?)

OpenStudy (anonymous):

cool deal! So lets go ahead and make the 3^2 into 9, and that -3(2)(-4) into +24 So youve got -3 +- sqrt(23) all over 4. I think that's about as far as you can take it, considering that 23 is prime

OpenStudy (anonymous):

Do you know how to put that into your calc to get an estimate?

OpenStudy (anonymous):

I don't know I will try

OpenStudy (anonymous):

how do you do the plus or minus on it ?

OpenStudy (anonymous):

You just have to do it twice, once with plus and once with minues

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ok I got 2.74456

OpenStudy (anonymous):

for the + or the -?

OpenStudy (anonymous):

so basically, when you are solving a quadratic equation, you are finding the x values for which y = 0

OpenStudy (anonymous):

that is +

OpenStudy (anonymous):

ok then cool so you understand that quadratic equations have 2 solutions?

OpenStudy (anonymous):

do I do the sqrt of 23 or the whole equation?

OpenStudy (anonymous):

Then with the - I got -17.2336...

OpenStudy (anonymous):

the sqrt of 23 only

OpenStudy (anonymous):

oh did not know that thanks

OpenStudy (anonymous):

it should be (3/4) - sqrt(23)/4 dont forget the /2a on the bottom

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok this is what i got -3sqrt 23/4=-.6020, then -3 - sqrt23-4= -5.3979

OpenStudy (anonymous):

the first one is + sorry

OpenStudy (anonymous):

shouldn't the sqrt be 33?

OpenStudy (anonymous):

probably lol. I'm sleepy. Make sure you divide by 4? the - one you just did looks a lil low. I think im going to go to bed, I have an 8am class tommorrow

OpenStudy (anonymous):

ok good night thanks for helping me your awsome